Five years ago Hemingway Inc. issued 6,000 30-year bonds with par values of $1,000 at a coupon rate of 8%. The bonds are now selling to yield 5%. The company also has 15,000 shares of preferred stock outstanding that pay a dividend of $6.50 per share. These are currently selling to yield 13%. Its common stock is selling at $21, and 200,000 shares are outstanding. Assume that the coupon payments are semi-annual. Calculate Hemingway's market value based capital structure. Round the values to the nearest dollar and the weights to two decimal places. Round PVF and PVFA values in intermediate calculations to four decimal places. Do not round other intermediate calculations.

Component Value Capital Structure
Debt $   %
Preferred stock %
Equity %
Total Capital $   %
Answer #1

Face value=    1000      
Years = 30-5=   25      
time (n) semiannual periods =25*2=    50      
Semiannual yield = 5%/2=   0.025      
Semiannual Coupon = 1000*8%/2=    40      


Bond price formula = Coupon amount * (1 - (1/(1+r)^n)/r + face value/(1+r)^n          
40*(1-(1/(1+0.025)^50))/0.025+ 1000/(1+0.025)^50          
          
1425.4347          
Number of bond's=   6000
Market value of debt= 6000*1425.434675 = 8,552,608.20      
          
Preferred stock yield=   13%      
Dividend=   6.5      
Price of preferred stock = Dividend/yield rate          
6.5/13%          
50          
Number of shares=   15000      
Market value of preferred stock= 15000*50=   750000      
          
Common stock price=   21      
number of Shares   200000      
          
Market value of Equity = 200000*21=   4200000      
          
          
Capital structure weight = Component Value/Total capital          
          
Component   VALUE   Capital structure weight  
Debt   8,552,608.20   0.6334  
preferred stock   750000   0.0555  
Equity   4200000   0.3111  
          
Total capital   13,502,608.20   1.0000